OpenStudy (anonymous):

What type of triangle is formed by joining the points D(7, 3), E(8, 1), and F(4, -1)? equilateral triangle isosceles triangle right triangle acute scalene triangle obtuse scalene triangle

OpenStudy (anonymous):

you would draw it ou,t connect the dots, evaluate the lengths then use law of cosines i believe

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

a^2 = b^2+c^2-2bccosA (1) b^2 = a^2+c^2-2accosB (2) c^2 = a^2+b^2-2abcosC.

OpenStudy (anonymous):

those are the law of cosine

OpenStudy (anonymous):

@acxbox22 what you think

OpenStudy (anonymous):

http://mathworld.wolfram.com/LawofCosines.html

OpenStudy (anonymous):

you use the law of cosines to find the angles

OpenStudy (anonymous):

then you use the angles to determine which is what

OpenStudy (anonymous):

c

OpenStudy (anonymous):

eualaterals have 3 60 degree angles

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

look at the picture in the link a is side a A is angle A. ive given you everything you need. if my answer lacks quality. sorry i created an account 5 minutes ago

OpenStudy (anonymous):

@dan815 what you think

OpenStudy (anonymous):

so what is the answer

OpenStudy (anonymous):

no this is how to do it. which gives you a way to do it yourself without spending time on research. which is even better than the answer, buddy

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

can somebody correct me